pooja195 (pooja195):

(2/x-2)=(4/x-3)-(1/x+1)??

OpenStudy (anonymous):

2/x-2=3/x-4 (2-2x)/x=(3-4x)/x 2-2x=3-4x 2x=1 x=1/2

pooja195 (pooja195):

thnk you!